OpenStudy (anonymous):

Derivative of y=C(1)sin(3x+C(2)) is the answer to this 3C(1)cos(3x+C(2)) or is it 1/3C(1)cos(3x+C(2))???

hartnn (hartnn):

while you've differentiated this, u'd use chain rule to get d/dx(3x+C(2)) = 3 so, that would get multiplied to it, and not divided. and its 3C(1)cos(3x+C(2))

OpenStudy (anonymous):

thank you.

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

okay wait, just a quick question. I don't know why i am confusing myself. But if I do u=3x+C2 then du=3x+C2 dx.... I though you solve for dx and you have du/3x+C2 = dx But this is how you would write this correctly?? dy=C1sin(u) du

OpenStudy (anonymous):

and the du = 3x+C2 dx... and thats how you get the 3?? I think I was just doing it backwards and confusing myself wtith the notation.

hartnn (hartnn):

just a sec, if u=3x+C1 then du = 3dx because C1 is constant(when you differentiate w.r.t x) and derivative of constant =0

hartnn (hartnn):

so, if you've put u=3x+C2 in y=C(1)sin(3x+C(2)) du = 3 dx dy = c1 sin u du = 3c1 sin (3x+c2).

hartnn (hartnn):

dy = c1 sin u du = 3c1 sin (3x+c2).dx

OpenStudy (anonymous):

when you use U substitution for integration it is the same process right? and still take the derivative the same way?

hartnn (hartnn):

yes, absolutely :)

OpenStudy (anonymous):

okay. thanks again, much appreciated.

hartnn (hartnn):

welcome again ^_^

OpenStudy (anonymous):

integrating cos 2x, would give me 2sin2x right?

hartnn (hartnn):

no, see this \(\int \cos 2x dx \\ put \: 2x =u , dx =du/2 \\ \int \cos u du/2 =1/2 (\sin u)+c =(1/2)sin 2x+c\) got this ?

OpenStudy (anonymous):

okay, thats how i did it, but i guess i am getting the rules confused between derivative and integration when using the chain rule.

OpenStudy (anonymous):

anyways, thats again for the 3rd time.

hartnn (hartnn):

are you starting integration ?

hartnn (hartnn):

if yes, and when you get time, you can go through this : http://openstudy.com/users/hartnn#/updates/50960518e4b0d0275a3ccfba